### v2

# SAR Strikes Back: A New Hope for RSVQA

Abstract
Remote Sensing Visual Question Answering (RSVQA) is a task that automatically extracts information from satellite images. It then processes a question to predict the answer from the images in textual form, helping with the interpretation of the image.
While different methods have been proposed to extract information from optical images with different spectral bands and resolutions, only recently have some preliminary studies started exploring very high-resolution Synthetic Aperture Radar (SAR) data. These studies leverage SAR’s ability to capture electromagnetic information and operate in all atmospheric conditions. However, no research has compared the results obtained using SAR and optical imagery or explored methods to fuse the two modalities effectively.
This work investigates the integration of SAR images into the RSVQA task exploring the most effective way to combine them with optical images.
In our research, we carry out a study on different pipelines for the task of RSVQA taking into account information from both SAR and optical data. To this purpose, we also present a dataset that allows for the introduction of SAR images in the RSVQA framework.
We study two different pipelines for RSVQA to include SAR modality and introduce a dataset enabling SAR-based RSVQA. The first model is an End-to-End approach while the second is a two-stage framework. In the latter, relevant information is extracted from SAR, before being translated into natural language to be used in the second step which only relies on a language model to provide the answer.
Our results show that the second pipeline achieves strong performance using SAR alone, yielding an improvement of nearly 10% in overall accuracy compared to the first one. We then explore various types of fusion methods to use SAR and optical images together. A fusion at the decision level achieves the best results on the proposed dataset, with a final F1-micro score of 75.00% and an F1-average of 81.21% for classification, as well as an overall accuracy of 75.49% for VQA. We show that SAR data offers additional information when fused with the optical modality, particularly for questions related to specific land cover classes, such as water areas.
Index Terms:
Deep Learning, Remote Sensing, Visual Question Answering, Multi-Modality, Natural Language Processing
I Introduction
Public and private sector initiatives are facilitating access to a wide range of remote sensing images. A well-known example is the Sentinel satellite constellation launched in 2014 as part of the European Union’s Copernicus programme. The mission provides free access to a wide range of imagery, including optical and Synthetic Aperture Radar (SAR) images. The optical images acquired by Sentinel-2 deliver high-resolution information based on the reflected sunlight. In contrast, SAR images, captured with the Sentinel-1 satellites, use radar signals which penetrate clouds and operate effectively at night. They provide data on surface roughness, moisture content, and other physical properties through radar backscatter. Sentinel-1 images contain two polarization channels, VV and VH, offering a more comprehensive representation of the scene.
The data coming from satellite images is used by scientists for a wide range of applications including environmental protection [1], biodiversity estimation [2] and demographic studies [3]. This data is also used by the public or journalists to identify events, conflicts, or the climate crisis [4]. However, it is time-consuming to extract information from remote sensing images. This work is performed by experts and often involves manual work, which can be a limiting factor considering the growth of data volumes. In addition, the extraction of information from satellite imagery is often limited to optical sensors operating in the visible spectrum, as they are easier to interpret. Indeed, interpreting SAR data is challenging due to geometric variability (with phenomena such as shortening, layout, and shadowing) and SAR image statistics called speckle.
To facilitate the extraction of information from remote sensing data, authors of [5] have proposed a new task in which the objective is to provide an answer to an open-ended question, expressed in natural language, about remote sensing images. This task is known as Remote Sensing Visual Question Answering (RSVQA) and originates from Visual Question Answering (VQA) [6]. In [5], the authors provide two datasets and a method that separately extracts textual and visual features from the questions and images. These features are then combined and passed to a multi-layer perceptron to choose the most appropriate answer to the question among a set of pre-defined ones.
Although this method shows promising results with optical images, it is not interpretable. In Prompt-RSVQA [7], this issue is addressed by dividing the process into two phases. In the first phase, the model identifies relevant information in the images to answer the questions. In the second phase, features are extracted from the question and combined with the class names identified in the images using a language model. This two-step method allows for a more detailed study of the classes detected in the images and a better understanding of potential prediction errors.
Both methods rely exclusively on optical images, much like the majority of the current state-of-the-art in RSVQA, with the exception of [8]. This trend is also prevalent across other tasks. The reliance on optical images stems from the greater complexity of SAR images, which has limited their usage in deep-learning based methodologies [9].
However, a fusion of SAR with optical images can leverage the unique strengths of each type of imagery. SAR images provide complementary information, such as detailed texture and surface characteristics. In addition, the content of optical images may not be visible at all due to atmospheric conditions or poor lighting. Integrating SAR with optical images can enhance overall analysis and interpretation by offering a comprehensive view of the scene. This combined approach has been shown to improve various applications, including object detection and land classification  [10].
In this work, we explore the use of SAR imagery in two RSVQA pipelines, End-to-End RSVQA and Prompt-RSVQA. We also study the effects of different fusion methods to combine optical and SAR imagery for RSVQA.
The paper is structured as follows: Section II reviews RSVQA, deep learning with SAR, and data fusion. Section III presents our pipelines and fusion methods, followed by our dataset in Section IV. Section V details evaluation metrics, with experiments and results in Section VI. We discuss findings in Section VII and conclude in Section VIII with insights on SAR in RSVQA and future directions.
II Related Works
RSVQA has been first introduced in [5]. Since then, many methods have been proposed, including approaches for identifying the best language models [11], fusion method between text and images [12], using segmentation maps to guide the fusion [13], and using text-based data augmentation [14]. The applications of RSVQA have broadened, with models now capable of providing diverse answers. For example, [15] explores an object detection model based on textual questions, while [16] introduces a model that answers questions requiring segmentation maps, vector maps, and object counting through a VQA interface, showcasing the versatility of RSVQA.
Research in RSVQA has predominantly focused on optical images, with SAR images being introduced only recently. The first study to incorporate SAR images [8], addressed land classification. Subsequent works explored questions related to scattering patterns and backscattering mechanisms [17], boat detection [18], and comparisons between the results got using optical images, obscured by clouds or captured at night, and SAR’s one [19].
Recent efforts in the remote sensing community have explored deep learning with SAR data. Hence, works on object detection [20], despeckling [21], volcano deformation detection [22] and images generation [23] have been proposed. In general, all techniques exclusively applied to optical imaging are now also extended to SAR imaging [9].
The interaction between SAR images and text is still in its early stages, primarily due to the inherent complexity of SAR data. While deep learning has shown limited effectiveness in tasks such as describing target sizes or counting objects [24], it performs better when identifying the relative positions of targets—such as indicating proximity to other objects—or providing density-based descriptions like ”few,” ”many,” or ”a lot.” Additionally, text has been used in conjunction with SAR data to generate synthetic images, helping to mitigate dataset imbalances and address the challenge of limited training samples [25].
One of the reasons for the predominant use of optical images is that they represent the visible spectrum. This makes them ideal for tasks such as object recognition [26], though they are limited by weather and lighting conditions. In contrast, SAR images, generated using radar signals, provide structural information and operate effectively in all weather and lighting conditions but are harder to interpret due to their complex patterns. Combining the optical and SAR modalities can leverage the strengths of both modalities, with optical images providing visual information and SAR images offering structural insights, resulting in a more comprehensive and accurate representation of the scene [27].
Various fusion techniques have been explored to combine optical and SAR images in deep learning pipelines. Early fusion, also known as raw data fusion, merges optical and SAR images before feature extraction, and has been applied in pansharpening [28] and Digital Surface Model (DSM) generation from multiangular images [29]. This approach is especially useful when the images are captured under similar conditions, such as using the same sensor or when the acquisition times are closely aligned. Halfway fusion, or feature-level fusion, extracts features from both modalities separately and then combines them for tasks like classification and change detection [30], offering a richer set of characteristics. Late fusion, or decision-level fusion, operates at the highest semantic level by merging the outputs of single-modality pipelines and is particularly effective in classification tasks [31]. The choice between halfway and late fusion, however, remains task-dependent. In classification, it is challenging to determine which technique performs better, as studies report that halfway fusion sometimes yields superior results [32, 33], while late fusion excels in other cases [34]. As noted in [35], the performance of either method is closely tied to the specifics of the problem and dataset.
Although some recent works have explored RSVQA with SAR images, all focus on very high-resolution SAR images [17] [18] [19]. This makes it difficult to compare the results with those obtained from optical images captured in clear weather and daylight conditions, especially when the questions differ for SAR and optical images and the images do not represent the same areas [36].
While it is known that SAR images perform better at night and on cloudy days, there has been no study in the RSVQA field on whether they provide unique information that optical images do not. Our work aims to address this gap by building on the findings of [8], examining the performance of two RSVQA pipelines when using SAR images alone or in combination with optical data.
III Method
Figure 1: The two pipelines proposed in this work. In both pipelines, we extract the visual information (an abstract visual feature vector $f_{i}$ for End-to-End RSVQA or a classification vector $f_{c}$ for Prompt-RSVQA) from one (i.e. $i=i_{o}$ or $i=i_{s}$) or two modalities (i.e. $i=(i_{o},i_{s})$). The End-to-End RSVQA pipeline processes both modalities separately. In this pipeline, we extract a feature vector $f_{q}$ from the question $q$. Both feature vectors are merged in a multi-modal vector $f_{a}$. In the Prompt-RSVQA pipeline, we convert the multi-label classification vector $f_{c}$ to text through a thresholding operation. This text is passed along the question to obtain a feature vector $f_{a}$ representing both the images and the question. In both pipelines, the multi-modal feature vector $f_{a}$ is used to predict the most likely answer $a$ from a set of possible answers $\mathcal{A}$.
The RSVQA task aims at providing an answer $a$ to a question $q$ from an image $i$. As seen in Section II, recent works on RSVQA only focused on the case in which $i=i_{o}$ is an optical image (with multiple spectral channels $n_{o}$). In this work, we introduce a SAR image $i_{s}$ (with multiple polarization channels $n_{s}$) to the problem setting. Hence, we formulate our task as providing an answer $a$ to a question $q$ from either a single modality (i.e. $i=i_{s}$ or $i=i_{o}$) or a pair of optical and SAR images $i=(i_{o},i_{s})$. Similarly to other works, we frame the RSVQA task as a classification problem, where the answer $a$ is predicted among a set of pre-defined answers $\mathcal{A}$.
To predict the right answer, we propose two methods shown in Figure 1: the first one is an End-to-End method (End-to-End RSVQA, described in Sub-section III-A). In this method, abstract features are extracted from the images $i$, as well as from the question $q$. These features are then merged to predict the answer $a$. The second one (Prompt-RSVQA, described in Sub-section III-B) is inspired by [7]. In this method, we first extract semantic information from the images. This is done through a separately trained multi-label classification network. This semantic information is then passed, along with the question $q$ to a language model to predict $a$. In this work, we also experiment with different methods for the extraction of the visual feature vector, with the proposition of three fusion mechanisms. The visual feature extractors are presented in Sub-section III-C.
III-A End-to-End RSVQA
In the End-to-End RSVQA pipeline, both the images $i$ and the question $q$ are processed separately. The different encoders used to obtain $f_{i}$, the feature vector of the images $i$, are described in section III-C. To obtain $f_{q}$, the feature vector of the question, we use a Recurrent Neural Network (RNN). We add a fully connected layer to map both vectors to new vectors $f^{\prime}_{i}$ and $f^{\prime}_{q}$ both of dimension $n_{a}$. Both of these vectors are merged into a new vector, $f_{a}$ through a point-wise multiplication. While the point-wise multiplication is a simple and fixed operation, the fully connected layers applied on $f_{i}$ and $f_{q}$ leave a degree of freedom for the network to reorganize the information in a way that helps the fusion of both vectors. Finally, the vector $f_{a}$, representing both the visual and textual information, is used as an input to a multi-layer perceptron to predict the final answer $a$. This MLP outputs a vector $y$, of size $|\mathcal{A}|$, giving a score for each possible answer from $\mathcal{A}$.
III-B Prompt-RSVQA
Prompt-RSVQA is organized into two stages. The first stage aims at extracting relevant semantic information from the image $i$. After obtaining $f_{i}$ (see Sub-section III-C), we predict $f_{c}$ a vector that represents classification scores through a multi-layer perceptron. More specifically, after applying a threshold on the prediction vector $f_{c}$, we obtain the set of classes describing the content of the image, which are concatenated in a text-based list called context $c$. Because this operation is non-differentiable, this first part of the pipeline is trained separately.
The second stage of Prompt-RSVQA takes as an input the question $q$ and the context $c$ extracted from $i$ in the first phase. Both of these texts are concatenated and fed into a transformer-based language model. To keep a similar setting to the pipeline of End-to-End RSVQA, we only use encoder layers and extract the feature vector $f_{a}$. This vector represents the information of the question and the context extracted from the image. Similarly, we use a multi-layer perceptron to predict the vector $y$ from $f_{a}$.
Figure 2: The different visual models studied in this work. (a) and (b) are mono-modalities models. In both cases, we use a CNN to extract abstract features $f_{i}$ (which can be used in End-to-End RSVQA) or high-level information $f_{c}$ (in our case, classes obtained after a MLP applied on the output of the convolutional layers, which can be used in Prompt-RSVQA). Similarly, the early fusion mechanism (c) uses a single CNN. However, in this setting, both $i_{o}$ and $i_{s}$ are stacked at the input of the CNN. In the halfway fusion (d), we concatenate the output of the convolutional layers before doing the classification. Finally, the late fusion mechanism adds a MLP to the classification scores obtained from each modality separately.
III-C Visual models
In this work, we examine different strategies to obtain a representation of the information (either abstract, $f_{i}$ for End-to-End RSVQA, or a classification $f_{c}$ in the case of Prompt-RSVQA) contained in the images $(i_{o},i_{s})$. To this end, we propose to examine five methods presented in Figure 2: optical only, SAR only, early fusion, halfway fusion and late fusion. In the mono-modal strategies (i.e. optical only or SAR only), we use a CNN as one would for classification. First, a series of convolutional layers are applied to obtain the visual feature vector $f_{i}$. This vector is then passed to a MLP which outputs a prediction vector $f_{c}$ indicating the scores for each possible class.
We first propose an early fusion method (Figure 2(c)). In this approach, the two modalities $i_{o}$ and $i_{s}$, with $n_{o}$ and $n_{s}$ channels respectively are stacked to create a single image with $n_{o}+n_{s}$ channels, which is then fed into the model. Similarly to the mono-modal strategies, we then use a CNN to extract the visual feature vector $f_{i}$ and a MLP to obtain the classification scores $f_{c}$.
The second fusion method, halfway fusion (Figure 2(d)), processes the two modalities separately through different CNNs. Before the final step, their feature maps are concatenated into a single feature map $f_{i}$, then passed to an MLP to obtain $f_{c}$.
The final method is named late fusion and is presented in Figure 2(e). With this strategy, the two modalities are processed separately through different CNNs. A vector representing classification scores is produced for each modality. These two vectors are then concatenated and passed to a MLP which finally produces the final decision vector, $f_{c}$.
IV Data
This section introduces a new dataset for multi-modal RSVQA: RSVQAxBEN-MM. This dataset is derived from three other datasets: BEN and BEN-MM, discussed in Sub-section IV-A and RSVQAxBEN, discussed in Sub-section IV-B. A sample of each of these datasets is shown in Figure 3.
Figure 3: Workflow from the initial BEN dataset containing the Sentinel-2 optical images and their CLC classes, to BEN-MM-61 where the respective Sentinel-1 images were added. We then move on to RSVQAxBEN in which questions and answers related to the optical images have been proposed. Finally, the dataset we propose in which such questions and answers are linked to both optical and SAR images.
IV-A Land-cover datasets: BEN and BEN-MM
BigEarthNet [37] (BEN) dataset is composed of 590’326 patches of Sentinel-2 images with 12 channels, acquired from over 10 European countries. Each patch is matched with the classes corresponding to the CORINE Land Cover (CLC) map of 2018. In the CLC nomenclature, the land cover classes are presented on three increasingly specific hierarchical levels. Level L1 introduces 5 more generic classes (e.g. Agricultural areas, Water bodies). Level L2 introduces 15 sub-classes (e.g. Arable land, Inland water). Finally, level L3 presents 44 classes at the finest level (e.g. Rice fields, Water courses). Over the three levels of information, the CLC nomenclature introduces 64 classes. In BEN, some classes have been deleted and others have been fused reaching a total of 19 classes. The train (60% of the patches), validation (20% of the patches) and test sets (20% of the patches) are defined randomly.
An extension of BEN, called BigEarthNet-Multi Modality dataset [38] (BEN-MM), has been later released adding to each Sentinel-2 patch of the original dataset the corresponding Sentinel-1 image, in the two polarisation channels (VV and VH) in dB. In this work, we also use another version of the BEN-MM dataset, with 61 classes as labels (see Section IV-B). In the rest of this manuscript, we refer to this dataset as BEN-MM-61.
IV-B RSVQA dataset: RSVQAxBEN
The RSVQAxBEN dataset is derived from BEN. Each of the 590,326 Sentinel-2 RGB patches is paired with 25 question/answer pairs, totaling 14,758,150 image/question/answer triplets.
The questions are based on each image’s CLC labels. As opposed to BEN, the 64 original classes of the CLC nomenclature are kept. However, two classes with the same name at different hierarchy levels (water bodies and pastures) are counted as a single label and the category Glaciers and perpetual snow is removed, leading to a total of 61 classes.
Questions fall into two categories: yes/no questions, in which the answer is ’yes’ or ’no’ (e.g. ’Is there a forest or water in this image?’) and land cover questions in which the answer is a list of classes (e.g. ’Besides forest, what classes are present in the image?’). The dataset is heavily unbalanced, with 80.7% of the questions being a yes/no question. In yes/no questions, conjunctions such as ’and’ and ’or’, adding a difficulty level to the question, are present in 72.3% of the cases. In 27.1% of the questions, two conjunctions are found. As opposed to BEN, the dataset split is divided according to the spatial location of the image: The westernmost patches are placed in the training set (66% of the dataset); the easternmost patches are put in the test set (23% of the dataset); finally, the other patches, representing 11% of the dataset, are put in the validation set. This choice allows the models to avoid biases due to geographical location. This split minimizes location-based biases but introduces a geographical generalization challenge, as some classes vary visually between Western and Eastern Europe.
IV-C Proposed dataset: RSVQAxBEN-MM
This work introduces the RSVQAxBEN-MM dataset. This dataset adds the SAR images from BEN-MM to the question/answer of RSVQAxBEN. An example is shown in Figure 3(d). Because it is derived from RSVQAxBEN, this dataset keeps the same choices, in terms of classes of interest and train / validation / test sets distribution. Each of the images / question / answer triplets is composed of the Sentinel-2 RGB patch as in RSVQAxBEN and adds the corresponding Sentinel-1 patch. The provided Sentinel-1 patches are composed of the VV, and VH channels in dB and the ratio between the two, providing $n_{s}=3$ SAR channels.
We use three channels, as combining VV and VH polarizations improves the separability of vegetation and artificial structures [39], while the VV/VH ratio effectively distinguishes volume from surface scattering [8, 40]. Both enhancing land cover classification in RSVQA tasks. In this context, with images expressed in dB scale, the ratio channel is calculated as the difference between VV and VH, normalized between 0 and 1.
The RGB representation of the SAR image is obtained by assigning VV to the R channel, VH to the green channel and the ratio to the blue channel.
IV-D Dataset evaluation)
While RSVQAxBEN-MM is the first dataset proposing SAR and optical modalities for RSVQA, it presents some limitations discussed in this section. The first limitation concerns the distribution of the CLC classes. Across the different patches, the 61 classes are strongly imbalanced. Among the three levels, the six most represented classes (Agricultural areas and Forest and Seminatural areas at L1, Arable land, Heterogeneous agricultural areas and Forests at L2, and Non-irrigated arable land at L3) together account for 54% of the 4’457’279 class occurrences in BEN-MM-61. This imbalance makes the classification task harder. In particular, it has been shown that since the questions and answers are based on the classes in each image, the use of weighted-losses during the training brings no benefit in VQA [8].
The second limitation concerns the biases present in the answers. To evaluate the biases in the dataset some scores are introduced in [41], namely: Uniform distribution, Prior distribution, and $L_{B_{score}}$. The uniform distribution is calculated as the inverse of the number of unique answers $A_{\text{unique}}$:
$\text{Uniform}=\frac{1}{A_{\text{unique}}}\quad.$ (1)
The Prior distribution is calculated as the ratio between the number of samples with the most common answer $A_{\text{common}}$ and the total number of samples ($N$):
$\text{Prior}=\frac{A_{\text{common}}}{N}\quad.$ (2)
Finally, assuming that Uniform does not equal 1 in a realistic scenario, the $L_{B_{\text{score}}}$ can be calculated:
$L_{B_{\text{score}}}=\frac{\text{Prior}-\text{Uniform}}{1-\text{Uniform}}\quad.$ (3)
All these scores work in a range of values between 0 and 1. Ideally, a perfect RSVQA dataset would have ${A_{unique}}=\dfrac{N}{A_{common}}$, which means that each answer has the same number of occurrences. In this perfect scenario, the $L_{B_{score}}$ would be equal to 0. In Table I the scores are applied to RSVQAxBEN-MM. When the question type is All, it measures the bias by treating the entire dataset as a single group, calculating how much the model favours the most common answer across all questions combined. These scores are also computed by question category, which are then combined on the whole dataset.
We can observe in All and Land Cover in Table I, that the Uniform distribution has a very low value, which means that there are many different types of answers. Instead in Yes/No the Uniform distribution has a value of 0.50 since there are only two types of answers.
Prior in All has a value of 0.52, which means that more than half of the answers are ‘no’ and is a direct result of the fact that in Yes/No as many as 63% of the answers are ‘no’. The dataset appears to be biased towards the most common answers, particularly when the dataset is viewed as a whole.
Question Type #samples #answers Most common Prior Uniform $\mathbf{L_{B_{score}}}$
All 2’953’125 26’875 no 0.52 0.00004 0.52
Land cover 529’413 26’873 None 0.13 0.00004 0.13
Yes/No 2’423’712 2 no 0.63 0.50 0.26
TABLE I: Analysis of RSVQAxBEN test set done in [41] using the Prior, Uniform and $L_{B_{score}}$ scores.
Despite these limitations, it is important to note that the distribution of classes and therefore answers is naturally unbalanced in reality as well.
V Performance evaluation
To evaluate our work, we use a VQA metric described in subsection V-C. In addition, we introduce metrics for classification in subsection V-A and subsection V-B that we use for the evaluation in the context prediction for Prompt-RSVQA.
V-A Classification
F1 Score: is defined as the harmonic mean of the precision (P) and recall (R), where an F1 score reaches its best value at 1 and worst score at 0:
$\textrm{F1}=2\cdot\frac{P\cdot R}{P+R},$ (4)
with
$P=\frac{TP}{TP+FP}\quad R=\frac{TP}{TP+FN},$ (5)
and $TP$ the number of true positive predictions, $FP$ the number of false positives and $FN$ the number of false negatives.
The F1 score is computed for each class. In the following, we report the F1-micro score and the F1-average score. The first one counts the total true positives, false negatives and false positives. While the latter is computed as a weighted arithmetic mean (with a per-class weight corresponding to the number of positives of the class) of the per-class F1.
V-B From classification to VQA
Match Ratio (MR): The MR computes the fraction of correctly classified samples, i.e. the samples whose predicted labels exactly correspond to the ground truth labels. This gives the following for Q samples:
$MR=\frac{1}{Q}\sum_{i=1}^{Q}I(f_{c_{i}}=\hat{f}_{c_{i}})\,$, (6)
where $f_{c_{i}}$ represents the actual labels (as a one-hot vector) for the $i_{th}$ sample, while $\hat{f}_{c_{i}}$ represents the labels predicted by the model for the same sample after thresholding. The identity function $I$ is 1 for an exact match and 0 otherwise.
Hamming Distance (HD): The HD is defined as the number of classes with a different prediction than the ground truth. It is defined for Q samples and N land cover categories as:
$HD=\frac{1}{Q}\sum_{i=1}^{Q}\sum_{j=1}^{N}I(f_{c_{{ij}}}=\hat{f}_{c_{ij}})\,$, (7)
where $f_{c_{{ij}}}$ and $\hat{f}_{c_{ij}}$ represent the prediction and ground truth of the $j^{th}$ land cover class for the $i^{th}$ sample.
V-C VQA
To evaluate VQA results, we define the percentage of correct answers as the accuracy. The global accuracy and per type of question (“Yes/No” or “Land cover” subsets) accuracy are provided.
VI Experiments
We evaluate our proposed method for RSVQA using optical images (Sentinel-2 RGB) and SAR images (VV and VH polarizations and their ratio). We first train image encoders on the BEN-MM-61 classification task and present the results in subsection VI-A. We compare this encoder with an ImageNet pre-trained encoder for the End-to-End RSVQA pipeline in subsection VI-B. Finally, we present our results obtained with the Prompt-RSVQA pipeline in subsection VI-C.
VI-A Classification
In the first classification experiments, optical and SAR images are used separately. Two networks of different depths, ResNet-50 and ResNet-152 are compared to assess whether greater depth yields better results. The last layer of the network is replaced with one that has the correct output number, i.e. 61. The sigmoid is used as an activation function. The performances of the single-modality encoders are presented in Table III (rows a,b,c,d). The performances are evaluated on the classification task using F1-micro and F1-average score. In addition, we assess the performances of the encoders for the RSVQA task using MR and HD.
Based on the single-modality results, we compare the three fusion methods introduced in subsection III-C (early fusion, halfway fusion and late fusion) with ResNet-50. The results are presented in the row e,f,g of Table  III.
In the early fusion, optical and SAR images are concatenated and inserted into a ResNet-50 pre-trained on ImageNet. The first layer of the ResNet is modified to have 6 channels as input instead of 3. In this layer, the weights are initialised using the weights resulting from the optical-only training of the ResNet-50 on BEN-MM-61 for the optical channels and the weights resulting from the SAR-only training of the ResNet-50 on BEN-MM-61 for the SAR channel.
In the halfway fusion, each modality is inserted in a frozen ResNet-50 pre-trained on the BEN-MM-61 classification task, with parameters set as in Table III[b,d]. The last layer of each network is deleted and the feature maps of the two modalities are concatenated. An average pooling, flattening, linear layer and sigmoid are applied to the concatenated feature maps to finally have the prediction of the classes. This last linear layer is retrained on the classification task with BEN-MM-61.
The late fusion takes as input the two optical and SAR images and inserts them into two different frozen ResNet-50 fully pre-trained on the classification task with the parameters set as in III[b,d]. The two decision vectors, before thresholding, are concatenated and fed into an MLP, which generates the final vector representing the detected classes in the images. Considering $n_{c}$ as the number of classes, the two vectors are concatenated after applying the sigmoid function, resulting in an input of size $n_{c}\times 2$ for the MLP. The MLP consists of three layers that progressively transform the vector size. The final output, has a dimention of 61, representing the probabilities of the predicted classes based on the combined information from both modalities. This MLP is also retrained on a classification task using BEN-MM-61.
We train all the models using Adam as an optimizer, a learning rate of $10^{-6}$, a batch size of 64 and a binary cross-entropy as the loss. The number of epochs used to train each model is displayed in Table III. In this work, all of the models are trained with an Nvidia V100 with 16GB GPU. The number of trainable hyperparameters per method is present in Table II.
Model Trainable Parameters
S2/S1 ResNet-152 58’268’797
S2/S1 ResNet-50 23’633’021
Early Fusion 23’642’429
Halfway Fusion 249’917
Late Fusion 67’405
TABLE II: Number of trainable parameters per model.
VI-B End-to-End RSVQA
In [5], the End-to-End RSVQA was tested only with optical data. In this study, we assess the robustness of the method using SAR images as inputs, and compare it to the results of optical only. For each modality, we compare ResNet-152 and ResNet-50 to extract features compatible with text. In both cases, an MLP composed of two layers is added at the end, after the application of a Hyperbolic Tangent function and a dropout of 0.5. First, the networks are frozen and their weights are initialized either with the pre-training on ImageNet or the classification task on BEN-MM-61 as described in section VI-A. In addition, for ResNet-50 initialized on ImageNet only, the network is unfrozen during the full End-to-End pipeline training. The results of the 10 experiments are presented in Table IV.
We trained all these models using Adam as an optimizer, a learning rate of $10^{-6}$, a batch size of 1024 and a cross-entropy as the loss for 20 epochs.
VI-C Prompt-RSVQA
In Prompt-RSVQA, the visual model extracts classes that are transformed in text to help the language model answer the question. Thus, the performance of the Prompt-RSVQA pipeline depends strongly on that of the visual model. DistilBERT is used as the LLM. After applying the Sigmoid activation function and a dropout of 0.5, two layers are added at the end to match the number of possible answers. The seven experiments conducted on the classification task are compared in the Prompt-RSVQA pipeline. Their settings are summarized in Table  V.
VII Results and discussion
VII-A Classification
By comparing the classification results for optical and SAR images in Table III[a,b] and Table III[c,d], we can see that the F1-micro score is 6 to 7% higher using optical images than using SAR images. This highlights the smaller discrimination capacity of SAR images, especially for classes such as green urban areas and wetlands, and in particular inland wetlands.
Table III[a,b,c,d] also shows that increasing the depth of the neural network, while keeping the same hyper-parameters and the same training time, does not lead to an improvement in classification. The performance of the ResNet-152 network is below the one of ResNet-50 for both modalities. In Table III[a,b] we can verify that optical results lose 1.6% Indeed, since our classification problem does not require such a complex representation, the additional depth of ResNet-152 may not provide significant benefits. On the contrary, it could worsen the results [42]. Moreover, the greater depth of ResNet-152 could limit its generalization capacities, decreasing its performance since the training set and our test set are geographically separated. This generalization issue could be greater in the case of optical data. Our results are in line with those obtained in [37], in which the shallowest network obtains the best results.
Modality Network Pre-trained on Epochs HD MR F1-micro F1-avg
a OPT ResNet-152 ImageNet 24 4.06 12.7% 73.00% 78.21%
b OPT ResNet-50 ImageNet 40 3.40 15.6% 74.60% 80.92%
c SAR ResNet-152 ImageNet 7 4.16 11.94% 67.23% 76.36%
d SAR ResNet-50 ImageNet 11 4.24 11.9% 67.26% 77.13%
e Early ResNet-50 BEN-MM/ImageNet* 31 4.03 13.7% 73.89% 80.54%
Fusion
f Halfway ResNet-50 BEN-$M$ 12 3.54 13.44% 74.10% 80.82%
Fusion
g Late ResNet-50 BEN-$M$ 2 3.35 14.3% 75.00% 81.21%
Fusion
TABLE III: Classification results on BEN-$M$ using ResNet-50/152, using a single modality as input (rows a-d) and both modalities (rows e-g).
* The network is pretrained on ImageNet, where the first layer is initialized using weights from BEN-MM models (rows b–d). See Section VI-A for details.
Then, we study the best type of fusion for our task considering only ResNet-50, as it performs the best for each modality individually. From Table III we can see that in terms of F1-micro, early fusion has the worst fusion results, followed by halfway fusion. Both have also slightly worse results than the optical image alone. Only late fusion does better than SAR and optical alone, performing the best.
The low results of early fusion are not surprising since it has been shown in [8] that inserting both SAR and optical images in one model may lead to primarily relying on optical channels. The fact that late fusion outperforms halfway fusion is not expected, since we have seen in Section II that the relative performances of each fusion depend on the nature of the task, its complexity, as well as on the dataset. Halfway fusion combines features extracted from each domain (optical and SAR) before classification. We believe that for our task, the features extracted from the two domains are not well aligned or containing redundant or conflicting information. The fusion may introduce noise rather than useful information. On the contrary, late fusion combines the predictions of the two modalities rather than their features. This may simplify the integration process since it operates on higher-level abstractions (e.g., class probabilities) rather than complex and potentially incompatible feature spaces. Interestingly, despite this simpler approach and having the fewest trainable parameters, as shown in Table II, the late fusion strategy still yields the best results among the three fusion schemes. Moreover, combining final predictions might reduce the impact of modality-specific noise. For instance, if one modality is particularly noisy or less informative for certain classes, its influence can be minimized when combined with the other modality’s more accurate predictions. Another advantage of the late fusion mechanism comes from the fact that both classifiers can be designed and trained separately, allowing to take advantage of specific strategies for each modality.
The BEN-MM-61 dataset is heavily unbalanced, with 6 classes that represent 54% of the overall dataset, namely: Agricultural areas, Forest and seminatural areas, Arable land, Heterogeneous agricultural areas, Forests, Non-irrigated arable land. Late fusion improves four of them compared to a model considering optical data only: Forest and seminatural areas, Arable land, Heterogeneous agricultural areas and Non-irrigated arable land. Halfway fusion improves on Forest and seminatural areas, Forests and Non-irrigated arable land. Finally, Early fusion shows an improvement on Agricultural areas only.
Figure 4: Percental variation in the behaviour of each L1 class based on the F1-average score, comparing results from the three fusion methods and SAR images alone to those from optical images.
Figure 4 shows the impact of each fusion method on individual L1 classes, reflected by changes in the F1-average scores. These scores account for dataset imbalance and compare the performance of three fusion methods and SAR images with optical images. The model applied to SAR images is generally less accurate for most L1 classes, with the exception of Water bodies. Halfway and late fusion make the most of the improvement of SAR images on the Water bodies, with even larger improvement with respect to SAR modality only. On the other hand, both these fusions are ineffective for the prediction of the Wetlands class, where only the early fusion can improve the prediction accuracy. The low performance of wetlands classification can be attributed to several factors. First, while SAR data, particularly HH polarization, is effective for classifying more structured areas like swamps and uplands, this polarization is not present in the dataset. Additionally, the class ”Swamps” is absent from our dataset, and many of the wetlands that are present, such as fens, bogs, and marshes, exhibit less structured vegetation, leading to misclassification [43]. Moreover, fusing SAR with optical data may introduce inconsistencies due to differences in resolution and spectral properties [44]. To enhance classification accuracy, the use of multispectral Sentinel-2 bands has been shown to improve detection, particularly for smaller wetlands [44]. Similarly, Digital Elevation Models (DEM) have proven useful in distinguishing wetland classes [43]. This aligns with the findings of [8], which demonstrated that adding a ratio channel, and thus incorporating more volumetric information, improved wetland classification by 2%.
In the Artificial area class, late fusion succeeds in improving only two L3 classes. In contrast, halfway fusion significantly improves the Artificial area class, particularly in three high-frequency classes, whereas early fusion benefits six smaller classes. In the case of Forest and seminatural areas, halfway fusion and late fusion improve almost the same classes, but late fusion has a major impact. Similarly, for Agricultural Areas both early and late fusion improve the performances, but late fusion improves more frequent classes. Our results show that different mergers improve different classes, but what remains common is that if it is not high-frequency classes, the overall performance does not improve significantly.
We next consider how these classification results may influence the VQA performance. When considering the Hamming distance, the best results are obtained with the late fusion. However, the best matching ratio remains the one obtained with optical images alone. This means that using the late fusion yields fewer errors on average. However, these errors are more spread in the different predictions compared to a model using optical data only.
VII-B End-to-End RSVQA
In our End-to-End RSVQA model applied to optical images, we notice that using a ResNet-50 instead of ResNet-152 yields an important performance drop, whether the weights are frozen or fine-tuned.
With SAR, performances remain low both when using ResNet-152 and ResNet-50. The strategy of using weights pre-trained on the BEN-MM classification task does not improve the results. This may be due to the misalignment between textual features and those extracted from the visual model.
Indeed, textual and visual representations come from different domains. This may cause a lack of direct correspondence between the two representations, complicating the alignment and fusion of information. This misalignment could be improved using more advanced fusion strategies [12].
Modality Network Pre-trained Frozen? Yes/No Land Cover Overall Overall
Accuracy Accuracy Accuracy $L_{B_{score}}$
a OPT ResNet-152 ImageNet Yes 80.03% 20.71% 69.94% 17.94%
b OPT ResNet-152 BEN-$M$ Yes 71.34% 13.44%- 61.49% 9.49%
c OPT ResNet-50 ImageNet Yes 72.17% 13.91% 62.26% 10,26%
d OPT ResNet-50 BEN-$M$ Yes 71.52% 13.86% 62.31% 10.31%
e OPT ResNet-50 ImageNet No 71.55% 12.17% 61.46% 9.49%
f SAR ResNet-152 ImageNet Yes 62.66% 15.23% 55.09% 3.09%
g SAR ResNet-152 BEN-$M$ Yes 61.97% 15.26% 54.51% 2.51%
h SAR ResNet-50 ImageNet Yes 71.54% 14.24% 61.79% 9.79%
i SAR ResNet-50 BEN-$M$ Yes 71.21% 13.50% 61.49% 9.49%
j SAR ResNet-50 ImageNet No 72.08% 14.14% 62.23% 10.23%
TABLE IV: End-to-End RSVQA results on the RSVQAxBEN-MM dataset.
VII-C Prompt-RSVQA
The results of the Prompt-RSVQA method build on the classification results presented in Section VII-A, even if improvements in classification performance do not directly translate to equivalent gains in VQA accuracy. Indeed, an important observation in classification is that the ResNet-50 model shows a 3.7% difference in MR between optical and SAR predictions. However, this results in a 3.62% increase in Accuracy.
For both SAR and optical images we can observe that the decrease in the F1-micro classification scores when increasing the network depth is reflected in the corresponding VQA results. It is interesting to notice in Table V[c,d] that with ResNet-152 the land cover accuracy obtained with SAR improves by 0.64%. However, these questions are under-represented in the dataset, which leads to an inferior overall accuracy compared to the one obtained with ResNet-50.
Regarding the fusion results, we observe a gradual improvement from early fusion, halfway fusion to late fusion. This is coherent with what we observe in the classification results. However, late fusion is the only one that outperforms a pipeline using optical images only, especially in improving the land cover accuracy. This is linked to the match ratio results.
Although these relatively simple fusion strategies already provide better performance, it should be noted that more complex approaches have also been explored. For instance, methods such as VisualBERT [45] have been applied in remote sensing VQA settings [46]. However, despite leveraging attention-based fusion mechanisms, the performance reported in that work remains below the level achieved by our simpler architecture.
Figure 6 represents the normalized confusion matrix for the Prompt-RSVQA method with a ResNet-50 backbone and a late fusion mechanism. On the left of Figure 6, the normalized confusion matrix with the 75 more frequent answers is presented ranked by frequency. On the right, two zooms of the confusion matrix are presented. Note that to enrich our interpretation, the zooms are presented in a non-normalized (and in log-scale) version. The top-right zoom highlights the 19 most frequent classes, while the bottom-right zoom displays the classes from the 55th to the 74th most frequent ones.
In the normalized confusion matrix, we observe that for the 30 most-frequent answers the trend is relatively correct, with a visible diagonal. In the non-normalised confusion matrix on the top right corner, the strong imbalance in the answers clearly appears. Indeed, 82% of questions of the test set are yes/no questions. As seen in Table V[g], our model can predict the answer to these questions with good accuracy. Another fact we can observe is that the model struggles to distinguish between classes that belong to the same branch of the CORINE Land Cover hierarchy. Two examples of this phenomenon can be illustrated in the blue bounding boxes with answers from the same hierarchy being ordered consecutively in the matrix. Visually, we show in Figure 5 some predictions made by Prompt-RSVQA on the test set.
Figure 5: Visual results from the test set. The two modalities, the prediction of OPTICAL and SAR image only and the late fusion prediction are represented.
Modality Network Pre-trained Frozen? Yes/No Land Cover Overall Overall
Accuracy Accuracy Accuracy $L_{B_{score}}$
a OPT ResNet-152 BEN-$M$ Yes 84.55% 24.93% 73.86% 21.86%
b OPT ResNet-50 BEN-$M$ Yes 86.07% 26.56% 75.40% 23.40%
c SAR ResNet-152 BEN-$M$ Yes 82.80% 21.28% 71.77% 19.77%
d SAR ResNet-50 BEN-$M$ Yes 82.94% 20.64% 71.78% 19.78%
e Early ResNet-50 BEN-$M$ Yes 85.54% 26.03% 74.88% 22.88%
Fusion
f Halfway ResNet-50 BEN-$M$ Yes 85.90% 25.91% 75.15% 23.15%
Fusion
g Late ResNet-50 BEN-$M$ Yes 86.07% 27.03% 75.49% 23.49%
Fusion
TABLE V: Prompt-RSVQA results on RSVQAxBEN-$M$ dataset using ResNet with different depth, modality as input and different types of data fusion.
Figure 6: Confusion matrices of the Prompt-RSVQA model with a ResNet-50 backbone and a late fusion mechanism. On the left-hand side is the normalized confusion matrix. On the right-hand side, two zooms in the logarithm scale are represented. ”CX” corresponds to class X in the CLC taxonomy, available online 1.
VIII Conclusion
In this article, we introduce a new dataset including SAR images as a modality for RSVQA: RSVQAxBEN-MM. To the best of our knowledge, our work represents the first attempt to explore RSVQA for land cover classification using high-resolution SAR imagery, compare its performance with optical images taken under clear conditions and during daylight, and investigate fusion strategies. Moreover, we extend two pipelines to use the SAR data as input. Our experiments with the End-to-End RSVQA model show that it only performs well with optical data when using ResNet-152 as an image encoder, indicating that it is not a robust model. We show that with Prompt-RSVQA, we can gain in performances with the SAR modality. This suggests that visual features may be too complex for effective interaction with textual features in networks that have not been extensively trained for SAR data. Furthermore, a model relying on SAR data alone does not achieve results as high as optical data. However, using late fusion, we notice that SAR can add relevant information, particularly in water-related classes. The choice of the fusion method, however, strongly depends on the dataset used. On our proposed dataset, which is imbalanced across classes, late fusion is the most effective method. This result may not hold for other datasets with a different data distribution. Indeed, while early and halfway fusion methods improve the classification results on more classes, these classes are under-represented in questions present in the dataset. As such, future work could be focused on developing new RSVQA datasets including SAR as a modality, with better question coverage and with improved data distributions. 11footnotetext: https://land.copernicus.eu/content/corine-land-cover-nomenclature-guidelines/html/